There are 10 students in a class: 6 boys and 4 girls. If the teacher picks a group of
3 at random, what is the probability that everyone in the group is a girl?
1
Expert's answer
2020-11-24T10:36:11-0500

Probability that nbr1 is a girl is p1 = 4/10.

Probability that nbr2 is a girl, given 3 girls left is 3/9

Probability that nbr3 is a girl given 2 girls left is 2/8

"P = (4*3*2)\/(10*9*8) = 1\/30."


Answer: 1/30.